OpenStudy (anonymous):

Josie rode her bike 18 miles at 12 mi-h and then rode another 24 miles at 15 mi-h . How long did she ride in all?

OpenStudy (igreen):

Take a printscreen..

OpenStudy (igreen):

18 / 12 = 1.5 hours 24 / 15 = 1.6 hours 1.5 + 1.6 = 3.1 hours. But take a printscreen of your question first, okay? Don't just put 3.1 hours as the answer.

OpenStudy (igreen):

@Rehamzia Are you there?

OpenStudy (anonymous):

Yes, thanks

OpenStudy (igreen):

Take a printscreen of your question.

OpenStudy (igreen):

Oh, okay. I thought it was a fill-in-the-blank question. You're good.

OpenStudy (igreen):

How many do you have left now?

OpenStudy (anonymous):

What is the answer?

OpenStudy (anonymous):

@iGreen What is the Final answer????

OpenStudy (igreen):

3.1 hours is the answer..

OpenStudy (anonymous):

OK thanks

OpenStudy (anonymous):

B. 3.1 because 18/12= 1.5 and 24/15 =1.3 add it up and you get 3.1 hrs